# DiffusionGPT[Uncaptioned image]: LLM-Driven Text-to-Image Generation System

### v1

@@ 4.3 Quantitative Results
The alignment between user preferences and the quantitative findings presented in Table 1 serves as strong evidence of the robustness and effectiveness of DiffusionGPT. To further evaluate the different generated results, we employed the aesthetic predictor and human feedback related reward model. By comparing the effects of our basic version with the baseline model SD1.5, the results in Table 1 demonstrate that our overall framework outperforms SD15 in terms of image-reward and aesthetic score, achieving improvements of 0.35% and 0.44% respectively.
4.4 Ablation Study
4.4.1 Tree-of-Thought and Human Feedback
@@
To validate the effectiveness of the designed components, we conducted a visual analysis of the performance achieved by incorporating different modules, as illustrated in Figure 8. In the figure, the variant labeled as “Random” represents the random sampling model. Notably, the random selection model generates a considerable number of images that do not align with the input prompts and lack semantic coherence. However, as we gradually incorporate the tree-of-thought (TOT) and human feedback (HF) modules into our system, a significant improvement in the quality of the generated images becomes evident. The images generated with the inclusion of TOT and HF modules exhibit enhanced realism, semantic alignment with the input prompts, and higher aesthetic appeal. This visual analysis showcases the advantage of our system in selecting superior models through the integration of TOT and HF components.
4.4.2 Prompt extension
To evaluate the effectiveness of the prompt extension agent, we conducted a comparison between the generation results using the original prompt and the extended prompt as inputs, as depicted in Figure 9. The extended prompts were designed to provide a richer and more detailed description of the desired image. Upon analysis, we observed that the extended prompts yielded notable improvements in the aesthetics and level of detail of the generated images. The inclusion of additional information in the extended prompts allowed for the generation of more visually appealing and artistically enhanced images.
4.5 User Study
To obtain real human preferences for the generated images, we conducted a user study comparing our model against a baseline model. We utilized image captions sourced from PartiPrompts[25], randomly selecting 100 prompts and generating four images for each prompt. Subsequently, we collected feedback from 20 users, who were asked to rate the superiority or equality of the images. This process resulted in approximately 400 votes for each base model (SD15 and SD XL). As depicted in Figure 7 and Figure 6, the results of the user study consistently revealed a clear preference for our model over the baseline. The users consistently expressed a distinct preference for the images generated by our model, indicating that they perceived them to be of higher quality or superiority compared to the baseline.
4.6 Limitations and Further works
Although DiffusionGPT has demonstrated the ability to generate high-quality images, there are still several limitations and our future plans are in the following ways:
Feedback-Driven Optimization We aim to incorporate feedback directly into the optimization process of LLM, enabling more refined pompt parse and model selection.
Expansion of Model Candidates To further enrich the model generation space and achieve more impressive results, we will expand the repertoire of available models.
Beyond Text-to-Image Tasks We intend to apply our insight to a broader set of tasks, including controllable generation, style migration, attribute editing, etc.
5 Conclusion
@@
We propose Diffusion-GPT, a one-for-all framework that seamlessly integrates superior generative models and efficiently parses diverse prompts. By leveraging Large Language Models (LLMs), Diffusion-GPT gains insights into the intent of input prompts and selects the most suitable model from a Tree-of-Thought (ToT) structure. This framework offers versatility and exceptional performance across different prompts and domains, while also incorporating human feedback through Advantage Databases. To sum up, Diffusion-GPT is training-free and can be easily integrated as a plug-and-play solution, offers an efficient and effective pathway for community development in this field.

### v2

@@ III-C Quantitative Results
The alignment between user preferences and the quantitative findings presented in Table II serves as strong evidence of the robustness and effectiveness of DiffusionAgent. To further evaluate the different generated results, we employed the aesthetic predictor and human feedback related reward model. By comparing the effects of our basic version with the baseline model SD1.5, the results in Table II demonstrate that our overall framework outperforms SD1.5 in terms of image-reward and aesthetic score, achieving improvements of 0.35% and 0.44% respectively.
III-D Ablation Study
III-D1 Tree-of-Thought and Human Feedback
@@
To validate the effectiveness of our designed components, we performed a visual analysis of various modules, as shown in Fig. 6. The “Random” variant represents the random sampling model, which produces many images that lack semantic coherence with the input prompts. In contrast, incorporating the tree-of-thought (TOT) and human feedback (HF) modules significantly enhances image quality, resulting in improved realism, semantic alignment, and aesthetic appeal. This analysis highlights the advantages of our system in selecting superior models through the integration of TOT and HF components.
III-D2 Prompt extension
To evaluate the effectiveness of the prompt extension agent, we compared generation results using original and extended prompts, as shown in Fig. 7. The extended prompts provided more comprehensive descriptions, enriching the context for generating visually appealing outputs. Our analysis revealed that extended prompts significantly improved the aesthetics and detail of the generated images, guiding the model to produce outputs with greater fidelity to the desired artistic style. Additionally, the detailed descriptions in the extended prompts enhanced the model’s ability to capture intricate nuances, resulting in images that exhibited higher intricacy and refinement compared to those generated with the original prompts.
III-E User Study
Figure 8: User Study: Comparing DiffusionAgent with SD1.5. Users show a strong preference for the expert models selected by DiffusionAgent across all 10 prompt categories.
Figure 9: Comparison of DiffusionAgent-Xl with base model.
To assess human preferences for generated images, we conducted a user study comparing our model to a baseline model using 100 randomly selected prompts from PartiPrompts[25], generating four images per prompt. Feedback was collected from 20 users who rated the images, resulting in approximately 400 votes for each baseline model (SD1.5 and SD XL). As shown in Fig. 8 and 9, the results consistently favored our model, with users perceiving its outputs as of superior quality versus the baseline.
IV Related Work
IV-A Text-based Image Generation
Initially, Generative Adversarial Networks (GANs) [4, 26] were widely used as the primary approach for Text-based image generation. However, the landscape of image generation has evolved, and diffusion models [5] have emerged as a dominant framework, especially when integrated with text encoders such as CLIP [12] and T5 [13], enabling precise text-conditioned image generation [14, 17, 9]. For instance, DAELL-2 [14] leverages CLIP’s image embeddings, derived from CLIP’s text embeddings through a prior model, to generate high-quality images. Similarly, Stable Diffusion [15] directly generates images from CLIP’s text embeddings. Imagen [17], on the other hand, utilizes a powerful language model like T5 [13] to encode text prompts, resulting in accurate image generation. To align text-to-image diffusion models with human preferences, recent methods [8, 2, 24] propose training diffusion models with reward signals. This ensures that the generated images not only meet quality benchmarks but also closely align with human intent and preferences.
IV-B Large Language Models (LLMs) for Vision-Language Tasks
The field of natural language processing (NLP) has witnessed a significant transformation with the emergence of large language models (LLMs) [10, 3, 21], which have demonstrated remarkable proficiency in human interaction through conversational interfaces. To further enhance the capabilities of LLMs, the Chain-of-Thought (CoT) framework [22, 28, 7, 29] has been introduced. This framework guides LLMs to generate answers step-by-step, aiming for superior final answers. Recent research has explored innovative approaches by integrating external tools or models with LLMs [18, 20, 23, 19]. For example, Toolformer [18] empowers LLMs with the ability to access external tools through API tags. Visual ChatGPT [23] and HuggingGPT [19] extend the capabilities of LLMs by enabling them to leverage other models to handle complex tasks that go beyond language boundaries. Drawing inspiration from these endeavors, we embrace the concept of LLMs as versatile tools and leverage this paradigm to guide T2I models to generate high-quality images.
V Conclusion
@@
We propose DiffusionAgent, a one-for-all framework that seamlessly integrates superior generative models and efficiently parses diverse prompts. By leveraging Large Language Models (LLMs), DiffusionAgent gains insights into the intent of input prompts and selects the most suitable model from a Tree-of-Thought (ToT) structure. This framework offers versatility and exceptional performance across different prompts and domains. To sum up, DiffusionAgent is training-free and can be easily integrated as a plug-and-play solution, offers an efficient and effective pathway for community development.